# Chapter 0 Hadron Production Processes

### v1

@@ Abstract
[Abstract]The experimental search for the pion – proposed in 1935 by Hideki Yukawa as the force carrier of the strong nucleon-nucleon interaction – was rewarded in 1947 when in cosmic ray photographic emulsion data a charged particle was identified with the proper mass of about 300 times the electron mass, completed three years later by the discovery of the neutral pion. Since then, accelerator-driven pion and meson (photo-)production on the nucleon and the associated production of new baryons have become the key elements for ground-breaking discoveries in numerous areas of particle and nuclear physics, from fundamental symmetries and their breaking to low-energy QCD dynamics, laying also foundations for modern elementary particle physics and the Standard Model. This article is an overview of eight decades of experimental and theoretical meson production physics, from isospin to charm and beyond, forming our understanding of hadrons and their interactions.
@@
Figure 1: Feynman diagram illustrating the production of a pion on a proton by an incident photon $\gamma$.
Nomenclature] AGS Alternating Gradient Synchrotron at BNL ANL Argonne National Laboratory ALICE A Large Ion Collider Experiment at CERN Bevatron Billions of eV synchrotron at LBL BEVALAC Bevatron plus HILAC linear accelerator at LBL BIC Bound States in the Continuum BNL Brookhaven National Laboratory BnGa Bonn-Gatchina Model BSE Bethe-Salpeter Equation CERN European Organization for Nuclear Research (Conseil Européen pour la Recherche Nucléaire) CC Model Coupled Channel Model CGLN amplitudes Chew, Goldberger, Low, and Nambu amplitudes CLAS CEBAF Large Acceptance Spectrometer at JLab c.m. Center-of-Mass Frame COSY Cooler Synchrotron CQM Constituent Quark Model ELSA Elektronen-Stretcher-Anlage (Electron-Stretcher-Facility) EM Electromagnetic FAIR Facility for Antiproton and Ion Research GANIL Grand Accélérateur National d’Ions Lourds GiM Giessen Model GlueX Gluonic Excitation Experiment GMO Gell-Mann-Okubo mass formula GRAAL Grenoble Accelerator LASer GSI Gesellschaft für Schwerionenforschung GW The George Washington University HADES High Acceptance DiElectron Spectrometer HILAC Heavy Ion Linear Accelerator at LBL ITEP Institute Theoretical and Experimental Physics JLab Jefferson Laboratory or Thomas Jefferson National Accelerator Facility (TJNAF) J-PARC Japan Proton Accelerator Research Complex LAMPF Los Alamos Meson Physics Facility LBL Lawrence Berkeley Laboratory LEPS Laser Electron Photon Experiment LET Low Energy Theorem LHC Large Hadron Collider at CERN LQCD Lattice Quantum Chromo-Dynamics MAMI Mainz Microtron OZI Rule Okubo–Zweig–Iizuka rule PDG Particle Data Group PSI Paul Scherrer Institute, previously SIN - Swiss Institute for Nuclear Research PWA Partial-Wave Analysis QCD Quantum Chromo-Dynamics QFT Quantum Field Theory QGP Quark–Gluon Plasma RICH Relativistic Heavy Ion Collider at BNL RIKEN National Research and Development Agency, Japan SAID Scattering Analyses Interactive Data SM Standard Model TRIUMF Tri-University Meson Facility VMD Model Vector Meson Dominance Model
@@
  • Short story of hadron physics.
@@ 1 Emergence of Subnuclear Particle Physics
The first recorded observation of particles different from the material forming the everyday-terrestrial environment was made as early as 1912 by Victor Franz Hess (1883-1964) in high-altitude balloon campaigns [2]. First considered merely as a curiosity, soon after the wide-reaching importance of Hess’ observations became evident. For his groundbreaking research, Hess was awarded the Nobel Prize in Physics in 1936, together with Carl David Anderson (1905-1991), who discovered the positron in cosmic rays [3], thereby opening the era of antimatter research. The modern understanding of air showers produced by cosmic rays is illustrated in Fig. 2.
@@
The epochal scientific revolution of quantum physics, speeding up in the decade after 1920, was necessary to make cosmic-ray research a flourishing enterprise. A big step forward was made by Patrick Maynard Stuart Blackett (1897-1974) by constructing in collaboration with Giuseppe Paolo Stanislao “Beppo” Occhialini (1907-1993) a coincidence Wilson cloud chamber. In 1933, he could confirm the positron and, in the years after, charged particle showers and pair production and pair annihilation processes, which he explained on the basis of the just formulated Dirac theory [^1] Paul Adrien Maurice Dirac (1902-1984) shared the 1933 Nobel Prize in Physics with Erwin Schrödinger (1887-1961) for the discovery of new productive forms of atomic theory. Cloud chambers, operated by supersaturated water vapor, were invented much earlier by the Scottish physicist Charles Thomson Rees Wilson (1869-1959), who was awarded the Nobel Prize in Physics in 1927 for his method of making the paths of electrically charged particles visible by condensation of vapor.
@@
In the late 1930s, Cecil Frank Powell (1903-1969) joined cosmic-ray research. He and Occhialini had optimized the technology and the use of photographic plates, which became essential for cosmic-ray research - and for many years after. Exposing the plates at high altitude to cosmic rays, Lattes, Muirhead, Occhialini, and Powell recorded for the first time the decay of a pi-meson, $\pi^{-}\to\mu^{-}+\bar{\nu}_{\mu}$, shown in Fig. 3 and only later understood to be caused by weak interaction [4, 5]. This measurement can be considered as the start of research on hadron production, first evidenced by the track recording the existence of the decaying particle and second by identifying that particle and its properties by the decay products. Basically, the same scheme is applied in modern production experiments at accelerator facilities, albeit on a much more involved technological level. In 1950, Powell and Occhialini won the Nobel Prize for their seminal work.
@@
In their first paper on the pion observation of 1947 [4], Lattes, Occhialini, and Powell write: We represent the primary mesons by the symbol $\pi$, and the secondary by $\mu$. In the second paper of 1947 [5], they call the particles casually pion and muon. There is good reason to state that this is the only documented birth and baptism certificate of the pion. The actual inventor of - at that time - such a hypothetical particle, however, was Hideki Yukawa (1907-1981) [6, 7] who postulated the existence of a meson (lent from the ancient Greek language for medium, i.e., between electron and nucleon masses). By that hypothesis, Yukawa could explain the short-ranged forces between nucleons in accordance with the extremely small size of nuclei. After the discovery of a medium-mass particle of about the required mass of about 200 $m_{e}$ in cosmic-ray showers, the new particle was initially considered as the Yukawa meson. By the work of Lattes et al., it became clear that the earlier found particle was not the Yukawa meson but the pion was the proper candidate while the muon was of a different nature. For a short while, the muon was in fact considered as the light meson but soon the distinct nature of the muon as a lepton and fermion [^2] Half-integer spin-particles are named after Enrico Fermi (1901-1954) who received the Noble price of 1938 for his demonstrations of the existence of new radioactive elements produced by neutron irradiation, and for his related discovery of nuclear reactions brought about by slow neutrons was understood.
@@
Interestingly, Yukawa himself apparently never claimed openly the name pion or pi-meson of his origin. In his original papers of 1935 and 1937, Yukawa denotes the new quantum as U-field [6, 7]. In his Nobel lecture of 1949 [8] and in a paper of that year [9] Yukawa seems to have accepted the naming of his meson as pion or pi-meson by using those names repeatedly.
@@
Figure 2: Illustration of an air shower resulting from the interaction of an incoming high energy cosmic particle with atmospheric nuclei. The figure is adapted from Ref. [10].
@@
In his own words, Yukawa summarizes in [9] the situation of 1949 by: This rather puzzling situation of the meson theory changed in 1947 due to the experiment by the Rome group on the decay of negative mesons on the one hand, and the discovery by the Bristol group of two kinds of mesons in cosmic rays on the other., where the Bristol group are Lattes et al. and the Rome group refers to an earlier paper of 1943 by Nereson and Rossi [11]. By denoting the muon as negative meson, Yukawa continues to use the notation which had been established after about 1935 when it was thought that there would be only a single meson.
@@
Figure 3: The photographic emulsion plate showing the first observation of - as known today - the weak decay of a pion, $\pi^{-}\to\mu^{-}+\bar{\nu}_{\mu}$. The muon-antineutrino $\bar{\nu}_{\mu}$ is not observed. Photo, including the original caption, taken from Ref. [4].
@@ 2 Hadron Production on Accelerators
As early as 1948, in parallel to cosmic ray research, new opportunities for hadron physics showed up when Gardner and Lattes published the first results on the Production of Mesons by the 184-Inch Berkeley Cyclotron [12], demonstrating that hadron research could be performed under laboratory conditions. The circulated beam of $380~{}\mathrm{MeV}$ alpha particles inside the cyclotron passed through the thin carbon target, producing mesons and a bunch of other particles. The negatively charged mesons were sorted out by the magnetic field and roughly focused on the edge of a stack of photographic plates. Besides using carbon targets, a few experiments were made on beryllium, copper, and uranium targets.
The modern era of particle physics - lasting until today - started with the advent of particle accelerators being able to deliver intense beams of highly energetic particles. In 1953, the Cosmotron synchrotron at BNL delivered proton beams with the design energy of $3.3~{}\mathrm{GeV}$. A year later, the Bevatron at Berkeley laboratory became operative with proton beams of even $6.2~{}\mathrm{GeV}$. With these machines, the cosmic ray mesons and V-particles could be produced and studied at the laboratory. The most eminent Bevatron result was - and is - the discovery of hadronic antimatter by observing first the antiproton in 1955 [13] and shortly after also the antineutron in 1956 [14], for which Owen Chamberlain (1920-1959) and Emilio Gino Segré (1905-1989) were honored with the Nobel Prize in Physics of 1959 for the discovery of the antiproton, a sub-atomic antiparticle.
@@ 1 Entering Strange Territory and Symmetry Violation
The years around 1947 actually played a key role in hadron discoveries. Rochester and Butler reported on Wilson cloud chamber observations indicating Evidence for the Existence of New Unstable Elementary Particles [15] showing an unusual decay pattern resembling a two-pronged fork, later denoted as V-particles [16], of mass $980~{}m_{e}\leq m<m_{p}$ ($m_{e}\simeq 511~{}\mathrm{keV}$ and $m_{p}\simeq 938.3~{}\mathrm{MeV}$ are the electron and proton rest masses, respectively) - fitting well with what shortly after was named the $K$-meson. In 1950, Hopper and Biswas of Melbourne observed another neutral particle decaying by the same V-like pattern seen by Rochester and Butler, but of a much heavier mass of about 2370 $m_{e}$ which matches surprisingly well the rest mass of the $\Lambda$ hyperon. The characteristic decay pattern of these strange particles led to the name V-particles [17]. Hence, the first hadrons - a meson and a baryon - with strangeness content were observed, although it still took another 15 years before a concise theoretical scheme was formulated.
@@
In the years after, the strangeness sector could be explored systematically, resulting in a whole spectrum of higher lying hyperon states. The observation of the non-conservation of parity (P) in hyperon decays was the first experimental datum of a broken symmetry at the level of elementary particles. In many cases, however, detailed spectroscopic studies were beyond the technological capabilities of that time. A typical case is the $\Lambda(1520)$ excitation. That $J^{\pi}=\frac{3}{2}^{-}$ excited state of the $\Lambda$ hyperon was observed originally as an elastic resonance in kaon scattering on a proton target in 1962 [18]. However, it lasted about 50 years and required an electroproduction experiment and the advanced detector setup at JLab before the pole position for the first hyperon of that rather sharp resonance ($\Gamma\simeq 18~{}\mathrm{MeV}$) could be determined precisely for $\Lambda(1520)$  [19].
@@
A far-reaching result of the Cosmotron and the Bevatron is related to parity non-conservation, i.e., left-right or mirror symmetry violation. Breaking mirror symmetry means that the image would be different from the original, e.g., placing pears in front of a symmetry-broken mirror might result in an image showing apples. The story started in 1949, Rosemary Brown of the Powell group [20, 21], identified on a emulsion plate, exposed to cosmic rays, an unexpected decay event. A hitherto unknown object, given the name k particle, with a mass of about $900~{}m_{e}$ decayed by two distinct theta and tau modes. For a long time, the validity of those data and a few rare follow-up observations by other groups were considered skeptically. By 1953, in total, 11 k particle events were known. When the Cosmotron and the Bevatron delivered their respective first beams, campaigns were started on that issue, rapidly increasing the number of k particle observations and the related theta- and tau-decays, now obtained under controllable laboratory conditions. The two separate decay modes were identified as given by two and three pions, respectively [22]. When Chinowsky and Steinberger [^3] Jack Steinberger (1921-2020) was a recipient of the 1988 Nobel Prize in Physics, along with Leon M. Lederman (1922-2018) and Melvin Schwartz (1932-2006), for the discovery of the muon neutrino) in 1954 could prove experimentally the intrinsic negative parity of pions [23], it was clear that the k particle suffers a parity-violating decay. Shortly after, the famous Wu experiment, performed by Chien-Shiung Wu (1912-1997), cleared skies by showing convincingly that weak interaction is indeed violating parity conservation [24] [^4] To the surprise of many Chien-Shiung Wu did NOT receive the Nobel prize for her pioneering Experimentum Crucis, thereby sharing the fate of many other women in science.. At that time, nomenclature was changed to $K$ meson or kaon. A posteriori, it’s evident that the puzzling $k$ particle decay was indeed the first observation of the weak decay of the kaon. It is noteworthy to mention that in 1954 Chinowsky and Steinberger also showed that charged and neutral pions have slightly different rest masses [25].
@@
Since their discovery, neutral kaons $K^{0}$ together with their anti-particle counterparts $\bar{K}^{0}$ have been a laboratory for research on fundamental aspects of particle physics. They were the first hadrons showing oscillation phenomena by the formation of short-lived ($K_{S}$, $t_{1/2}\approx 8\cdot 10^{-11}~{}\mathrm{s}$) and long-lived ($K_{L}$, $t_{1/2}\approx 5\cdot 10^{-8}~{}\mathrm{s}$) superpositions of the flavor (strangeness) eigenstates $K^{0}$, $S=-1$ and $\bar{K}^{0}$, $S=+1$, respectively. The differences in lifetime are due to the differences in the phase space available for the two-pion and the three-pion decay channels: the larger two-pion phase space causes a large decay width and shorter lifetime than the three-pion decay.
@@
As early as 1967, Andrey Dmitrievich Sakharov (1921-1989) [^5] In later years Sakharov engaged in human rights activities, turning into a Soviet dissident, and received in 1975 the Nobel prize for Peace for emphasizing human rights around the world. published in [31] the famous three Sakharov conditions on the cosmological matter-antimatter asymmetry problem. The second paradigm demands violation of C and CP symmetry. The modern understanding of CP violation on the microscopic level in hadrons is related to a CP-violating phase factor in the Cabibbo–Kobayashi–Maskawa (CKM) matrix [32] (named after Nicola Cabibbo (1935-2010), Makoto Kobayashi (born 1944), and Toshihide Maskawa (1940-2021))). Kobayashi and Maskawa were awarded half of the 2008 Nobel Prize in Physics for the discovery of the origin of the broken symmetry which predicts the existence of at least three families of quarks in nature. The CKM matrix describes the mixing of quarks under weak interactions, which is also relevant in another open nuclear physics problem, namely neutrinoless double beta decay. Taking together, these research areas are connecting sub-nuclear and cosmological phenomena. The present status about conserved symmetry is that time reversal combined with CP symmetry is conserved, forming together the triple-symmetry CPT.
As a side remark it might be of interest that Richard Feynman (1918-1988) is said to have placed at a conference in 1955 a 50 US Dollar bet against parity violation [33], underlining how unreal such a conjecture appeared at the time of the first rumors of an experimental signature. Feynman is eminently known for his contributions to the development of quantum electrodynamics for which he received the Nobel Prize in Physics in 1965 jointly with Julian Schwinger (1918-1994) and Shinichiro Tomonaga (1906-1979). Moreover, Feynman’s name is engraved forever in the Feynman diagrams which are the indispensable tools for modern QFT research in any theoretical context.

### v2

@@ Abstract
[Abstract]The experimental search for the pion – Hideki Yukawa’s meson proposed in 1935 as the force carrier of the strong nucleon-nucleon interaction – was rewarded in 1947 when in cosmic ray photographic emulsion data a charged particle was identified with the proper mass of about 300 times the electron mass, completed three years later by the discovery of the neutral pion. Since then, accelerator-driven pion and meson (photo-) production on the nucleon, along with the associated production of new baryons, has become a key element for ground-breaking discoveries in numerous areas of particle and nuclear physics, ranging from fundamental symmetries and their breaking to low-energy QCD dynamics to laying the foundations for modern elementary particle physics and the Standard Model. This article provides an overview of almost a century of experimental and theoretical research on meson production, from isospin to charm and beyond, shaping our understanding of hadrons, their spectra, structure, and interactions.
@@
Figure 1: Feynman diagram describing the production of a charged pion (dashed line) and a neutron (full line) by an incident photon $\gamma$ (wavy line), incident on a proton (full line) .
Nomenclature] AGS Alternating Gradient Synchrotron at BNL ANL Argonne National Laboratory ALICE A Large Ion Collider Experiment at CERN Bevatron Billions of eV synchrotron at LBL BEVALAC Bevatron plus HILAC linear accelerator at LBL BIC Bound States in the Continuum BNL Brookhaven National Laboratory BnGa Bonn-Gatchina Model BSE Bethe-Salpeter Equation CERN European Organization for Nuclear Research (Conseil Européen pour la Recherche Nucléaire) CC Coupled Channel CC Model Coupled Channel Model CBELSA Crystal-Barrel-Detector at ELSA CGLN amplitudes Chew, Goldberger, Low, and Nambu amplitudes CLAS CEBAF Large Acceptance Spectrometer at JLab c.m. Center-of-Mass Frame COSY Cooler Synchrotron CQM Constituent Quark Model CMS Compact Muon Solenoid ELSA Elektronen-Stretcher-Anlage (Electron-Stretcher-Facility) EM Electromagnetic FAIR Facility for Antiproton and Ion Research FNAL Fermi National Laboratory GANIL Grand Accélérateur National d’Ions Lourds GiM Giessen Model GlueX Gluonic Excitation Experiment GMO Gell-Mann-Okubo mass formula GRAAL Grenoble Accelerator LASer GSI Gesellschaft für Schwerionenforschung GW The George Washington University HADES High Acceptance DiElectron Spectrometer HILAC Heavy Ion Linear Accelerator at LBL ITEP Institute of Theoretical and Experimental Physics JINR Joint Institute for Nuclear Research JLab Jefferson Laboratory or Thomas Jefferson National Accelerator Facility (TJNAF) J-PARC Japan Proton Accelerator Research Complex LAMPF Los Alamos Meson Physics Facility LBL Lawrence Berkeley Laboratory LEPS Laser Electron Photon Experiment LET Low Energy Theorem LHC Large Hadron Collider at CERN LQCD Lattice Quantum Chromo-Dynamics MAMI Mainz Microtron OZI Rule Okubo–Zweig–Iizuka rule PDG Particle Data Group PSI Paul Scherrer Institute, previously SIN - Swiss Institute for Nuclear Research PWA Partial-Wave Analysis QCD Quantum Chromo-Dynamics QED Quantum Electrodynamics QFT Quantum Field Theory QGP Quark–Gluon Plasma RICH Relativistic Heavy Ion Collider at BNL RIKEN National Research and Development Agency, Japan SAID Scattering Analyses Interactive Data SIN see PSI SM Standard Model TAPS Two-Arm Photon Spectrometer TRIUMF Tri-University Meson Facility VMD Vector Meson Dominance Model
@@
  • A short story of hadron physics.
@@ 1 Emergence of Subnuclear Particle Physics
The first recorded observation of particles different from the material forming the everyday terrestrial environment was made as early as 1912 by Victor Franz Hess (1883-1964) during high-altitude balloon campaigns [2]. First considered merely a curiosity, the wide-reaching importance of Hess’s observations soon became evident. For his groundbreaking research, Hess was awarded the Nobel Prize in Physics in 1936, along with Carl David Anderson (1905-1991), who discovered the positron in cosmic rays [3], thereby marking the beginning of the era of antimatter research. The modern understanding of air showers produced by cosmic rays is illustrated in Fig. 2.
@@
The epochal scientific revolution of quantum physics, accelerating in the decade after 1920, served as a primer for Cosmic-ray research, which has made it a thriving field of study. A significant step forward was made by Patrick Maynard Stuart Blackett (1897-1974) in collaboration with Giuseppe Paolo Stanislao “Beppo” Occhialini (1907-1993) in constructing a coincidence Wilson cloud chamber. In 1933, he confirmed the positron and, in the years that followed, charged particle showers, pair production, and pair annihilation processes, which he explained based on the newly formulated Dirac theory [^1] Paul Adrien Maurice Dirac (1902-1984) shared the 1933 Nobel Prize in Physics with Erwin Schrödinger (1887-1961) for the discovery of new productive forms of atomic theory.. Cloud chambers, operated by supersaturated water vapor, were invented much earlier by the Scottish physicist Charles Thomson Rees Wilson (1869-1959), who was awarded the Nobel Prize in Physics in 1927 for his method of making the paths of electrically charged particles visible through the condensation of vapor.
@@
In the late 1930s, Cecil Frank Powell (1903-1969) joined cosmic-ray research. He and Occhialini optimized the technology and the use of photographic plates, which became essential for cosmic-ray research—and for many years thereafter. Exposing the plates at high altitude to cosmic rays, Lattes, Muirhead, Occhialini, and Powell recorded, for the first time, the decay of a pi-meson, $\pi^{-}\to\mu^{-}\bar{\nu}_{\mu}$, shown in Fig. 3 and only later understood to be caused by weak interaction [4, 5]. This measurement can be considered the start of research on hadron production, first evidenced by the track recording the existence of the decaying particle, and second by identifying that particle and its properties through the decay products. Basically, the same scheme is applied in modern production experiments at accelerator facilities, albeit at a much more involved technological level. In 1950, Powell and Occhialini won the Nobel Prize for their seminal work.
@@
In their first paper on the pion observation of 1947 [4], Lattes, Occhialini, and Powell write: We represent the primary mesons by the symbol $\pi$, and the secondary by $\mu$. In the second paper of 1947 [5], they casually refer to the particles as pions and muons. There is a good reason to state that this is the only documented birth and baptism certificate of the pion. The actual inventor of - at that time - such a hypothetical particle, however, was Hideki Yukawa (1907-1981) [6, 7] who postulated the existence of a meson (lent from the ancient Greek language for medium, i.e., between electron and nucleon masses). By that hypothesis, Yukawa could explain the short-range forces between nucleons in accordance with the tiny size of the nuclei. After the discovery of a medium-mass particle of about 200 $m_{e}$ in cosmic-ray showers, the new particle was initially considered to be the Yukawa meson. By the work of Lattes et al., it became clear that the earlier found particle was not the Yukawa meson, but the pion was the proper candidate, while the muon was of a different nature. For a short while, the muon was, in fact, considered as the light meson but soon the distinct nature of the muon as a lepton and fermion [^2] Half-integer spin-particles are named after Enrico Fermi (1901-1954) who received the Nobel Prize in 1938 for his demonstrations of the existence of new radioactive elements produced by neutron irradiation, and for his related discovery of nuclear reactions brought about by slow neutrons. was understood.
@@
Interestingly, Yukawa himself apparently never claimed the name pion or pi-meson openly as his origin. In his original papers of 1935 and 1937, Yukawa denotes the new quantum as U-field [6, 7]. In his Nobel lecture of 1949 [8] and in a paper from that year [9], Yukawa seems to have accepted the naming of his meson as pion or pi-meson by using those names repeatedly.
@@
Figure 2: Illustration of an air shower resulting from the interaction of an incoming high-energy cosmic particle with atmospheric nuclei. The figure is adapted from Ref. [10].
@@
In his own words, Yukawa summarizes in [9] the situation of 1949 by stating: This rather puzzling situation of the meson theory changed in 1947 due to the experiment by the Rome group on the decay of negative mesons on the one hand, and the discovery by the Bristol group of two kinds of mesons in cosmic rays on the other., where the Bristol group refers to Lattes et al. and the Rome group refers to an earlier paper from 1943 by Nereson and Rossi [11]. By denoting the muon as negative meson, Yukawa continues to use the notation that had been established after about 1935, when it was thought that there would be only a single meson.
@@
Figure 3: The photographic emulsion plate showing the first observation of - as known today - the weak decay of a pion, $\pi^{-}\to\mu^{-}\bar{\nu}_{\mu}$. The muon-antineutrino $\bar{\nu}_{\mu}$ is not observed. Photo, including the original caption, taken from Ref. [4].
@@ 2 Hadron Production on Accelerators
As early as 1948, in parallel with cosmic ray research, new opportunities for hadron physics emerged when Gardner and Lattes published the first results on the Production of Mesons by the 184-Inch Berkeley Cyclotron [12], demonstrating that hadron research could be performed under laboratory conditions. The circulating beam of $380~\mathrm{MeV}$ alpha particles inside the cyclotron passed through the thin carbon target, producing mesons and a bunch of other particles. The negatively charged mesons were sorted out by the magnetic field and roughly focused at the edge of a stack of photographic plates. Besides using carbon targets, a few experiments were conducted on beryllium, copper, and uranium targets.
The modern era of particle physics — lasting until today — started with the advent of particle accelerators that were able to deliver intense beams of highly energetic particles. In 1953, the Cosmotron synchrotron at BNL delivered proton beams with a design energy of $3.3~\mathrm{GeV}$. A year later, the Bevatron at Berkeley Laboratory became operational with proton beams of even $6.2~\mathrm{GeV}$. With these machines, cosmic ray mesons and V-particles could be produced and studied in the laboratory. The most eminent Bevatron result was — and is — the discovery of hadronic antimatter by first observing the antiproton in 1955 [13] and shortly after also the antineutron in 1956 [14], for which Owen Chamberlain (1920-1959) and Emilio Gino Segré (1905-1989) were honored with the Nobel Prize in Physics in 1959 for the discovery of the antiproton, a sub-atomic antiparticle.
@@ 1 Entering Strange Territory and Symmetry Violation
The years around 1947 actually played a key role in hadron discoveries. Rochester and Butler reported on Wilson cloud chamber observations indicating Evidence for the Existence of New Unstable Elementary Particles [15], showing an unusual decay pattern resembling a two-pronged fork, later denoted as V-particles [16], with a mass of $980~m_{e}\leq m<m_{p}$ ($m_{e}\simeq 511~\mathrm{keV}$ and $m_{p}\simeq 938.3~\mathrm{MeV}$ are the electron and proton rest masses, respectively) - fitting well with what was shortly thereafter named the $K$-meson. In 1950, Hopper and Biswas of Melbourne observed another neutral particle decaying by the same V-like pattern observed by Rochester and Butler, but with a much heavier mass of about 2370 $m_{e}$, which surprisingly matches the rest mass of the $\Lambda$ hyperon. The characteristic decay pattern of these strange particles, shown in Fig. 4, led to the name V-particles [17]. Hence, the first hadrons - a meson and a baryon - with strangeness content were observed, although it still took another 15 years before a concise theoretical scheme was formulated.
@@
In the years that followed, the strangeness sector could be explored systematically, resulting in a whole spectrum of higher-lying hyperon states. The observation of the non-conservation of parity (P) in hyperon decays was the first experimental datum of a broken symmetry at the level of elementary particles. In many cases, however, detailed spectroscopic studies were beyond the technological capabilities of that time. A typical case is the $\Lambda(1520)$ excitation. That $J^{\pi}=\frac{3}{2}^{-}$ excited state of the $\Lambda$ hyperon was originally observed as an elastic resonance in kaon scattering on a proton target in 1962 [18]. However, it lasted about 50 years and required an electroproduction experiment and the advanced detector setup at JLab before the pole position for the first hyperon of that rather sharp resonance ($\Gamma\simeq 18~\mathrm{MeV}$) could be determined precisely for $\Lambda(1520)$  [19].
@@
A far-reaching result of the Cosmotron and the Bevatron is related to parity non-conservation, i.e., left-right or mirror symmetry violation. Breaking mirror symmetry means that the image would be different from the original, e.g., placing pears in front of a symmetry-broken mirror might result in an image showing apples. The story started in 1949, Rosemary Brown of the Powell group [20, 21], identified on an emulsion plate, exposed to cosmic rays, an unexpected decay event. A hitherto unknown object, given the name $k$ particle, with a mass of about $900~\mathrm{m_{e}}$ decayed by two distinct theta ($\theta^{+}$) and tau ($\tau^{+}$) modes. For a long time, the validity of those data and a few rare follow-up observations by other groups were considered questionable. By 1953, a total of 11 $k$ particle events were known. When the Cosmotron and the Bevatron delivered their respective first beams, campaigns were initiated on that issue, rapidly increasing the number of $k$ particle observations and the related theta- and tau-decays, which are now obtained under controllable laboratory conditions. The two separate decay modes were identified as given by two and three pions, respectively [22]. When Chinowsky and Steinberger [^3] Jack Steinberger (1921-2020) was a recipient of the 1988 Nobel Prize in Physics, along with Leon M. Lederman (1922-2018) and Melvin Schwartz (1932-2006), for the discovery of the muon neutrino. in 1954 could prove experimentally the intrinsic negative parity of pions [23], it was clear that the $k$ particle suffers a parity-violating decay. Shortly after, the famous Wu experiment, performed by Chien-Shiung Wu (1912-1997), cleared the skies by showing convincingly that weak interaction indeed violates parity conservation [24] [^4] To the surprise of many, Chien-Shiung Wu did NOT receive the Nobel Prize for her pioneering Experimentum Crucis, thereby sharing the fate of many other women in science. At that time, the nomenclature was changed to $K$ meson or kaon. A posteriori, it’s evident that the puzzling $k$ particle decay was indeed the first observation of the weak decay of the kaon. It is noteworthy to mention that, in 1954, Chinowsky and Steinberger also showed that charged and neutral pions have slightly different rest masses [25].
Describing the striking discovery of parity non-invariance as a fundamental interaction, which is a game-changer for elementary particle physics, does not overemphasize the importance of that observation. In 1956, Tsung-Dao (T.D.) Lee (1926-2024) and Chen Ning (C.N.) Yang (1922-2025) formulated a concise theory explaining phenomena that, at that time, were totally unexpected and counterintuitive. In their seminal paper titled Question of Parity Conservation in Weak Interactions [26], honored by the 1957 Nobel Prize in physics, they wrote: The way out of the difficulty is to assume that parity is not strictly conserved, so that $\theta^{+}$ and $\tau^{+}$ are two different decay modes of the same particle, which necessarily has a single mass value and a single lifetime. The Nobel Prize was awarded in equal parts to C.N. Yang and T.D Lee for their penetrating investigation of the so-called parity laws, which have led to important discoveries regarding elementary particles.
@@
Since their discovery, neutral kaons $K^{0}$ together with their anti-particle counterparts $\bar{K}^{0}$ have been a laboratory for research on fundamental aspects of particle physics. They were the first hadrons to show oscillation phenomena through the formation of short-lived ($K_{S}$, $t_{1/2}\approx 8\cdot 10^{-11}~\mathrm{s}$) and long-lived ($K_{L}$, $t_{1/2}\approx 5\cdot 10^{-8}~\mathrm{s}$) superpositions of the flavor (strangeness) eigenstates $K^{0}$, S=1S=-1 and $\bar{K}^{0}$, $S=+1$, respectively. The differences in lifetime are due to the differences in the phase space available for the two-pion and three-pion decay channels: the larger two-pion phase space causes a larger decay width and a shorter lifetime than that of the three-pion decay.
@@
As early as 1967, Andrey Dmitrievich Sakharov (1921-1989) [^5] In later years, Sakharov engaged in human rights activities, turning into a Soviet dissident, and received in 1975 the Nobel Prize for Peace for emphasizing human rights around the world. published in [32] the famous three Sakharov conditions on the cosmological matter-antimatter asymmetry problem. The second paradigm demands a violation of CC and $CP$ symmetry. The modern understanding of CP violation at the microscopic level in hadrons is related to a CP-violating phase factor in the Cabibbo–Kobayashi–Maskawa (CKM) matrix [33] (named after Nicola Cabibbo (1935-2010), Makoto Kobayashi (born 1944), and Toshihide Maskawa (1940-2021)). Kobayashi and Maskawa were awarded half of the 2008 Nobel Prize in Physics for the discovery of the origin of broken symmetry, which predicts the existence of at least three families of quarks in nature. The CKM matrix describes the mixing of quarks under weak interactions, which is also relevant in another open nuclear physics problem, namely, neutrinoless double beta decay. In combination, these research areas connect sub-nuclear and cosmological phenomena. The present status is that time reversal, combined with $CP$ symmetry, is conserved, forming together the triple-symmetry $CPT$.
As a side remark, it might be of interest that Richard Feynman (1918-1988) is said to have placed a 50 US Dollar bet against parity violation [34] at a conference in 1955, underlining how unreal such a conjecture appeared at the time of the first rumors of an experimental signature. Feynman is eminently known for his contributions to the development of quantum electrodynamics, for which he received the Nobel Prize in Physics in 1965, jointly with Julian Schwinger (1918-1994) and Shinichiro Tomonaga (1906-1979). Moreover, Feynman’s name is engraved forever in Feynman diagrams, which are indispensable tools for modern QFT research in any theoretical context. The diagram shown in Fig. 1 summarizes the topic of this article in a well-defined, albeit rather condensed, graphical language of general rules [35, 36].